# MACEval: A Multi-Agent Continual Evaluation Network for Large Models

### v1

@@ 5 Conclusion
This work introduces MACEval, a novel dynamic evaluation framework for large generative models, which provides effective solutions to avoid data contamination and reduce human participation by a multi-agent system-based automatic evaluation network. To identify the upper bound of performance and evaluate continually, we propose an AUC-inspired metric, further incorporating the evaluation topologies to measure the overall performance of large models. Evaluation results on five high-profile capabilities, including visual perception, text comprehension, math, algorithm reasoning, and coding, demonstrate the effectiveness, efficiency, and flexibility of MACEval compared to existing benchmarks. We believe that employing large models as agents to form the autonomous evaluation network presents a promising direction for achieving safe and fair evaluation.
Appendix
Appendix A Limitations and Societal Impact
Although our study provides validation of the MACEval framework in various tasks, there are several potential limitations acknowledged below. First, our experiments are limited to 9 tasks across five common capabilities, encompassing a specific range of topics. Incorporating a broader spectrum of tasks, such as those within a specialty area, helps to yield more comprehensive insights into MACEval’s applicability. Second, we apply evaluation topologies with fewer hops, i.e., nodes and edges, to preliminarily demonstrate the effectiveness of the proposed framework. In the future, it is possible to construct larger evaluation networks to include a richer array of capability assessments. Third, in MACEval, the evaluation data stream consists of question-answer pairs graded by difficulty. One of the principles in constructing open-ended tasks is that the difficulty should be quantifiable, which imposes limitations on the implementation of many tasks.
Meanwhile, our work holds the potential for significant societal impact, both positive and negative. First, the evaluation is normally considered as important as training in developing large models. An insightful evaluation can offer promising directions to large model developers. The main motivation of this work is to build an automatic, scalable, and trustworthy evaluation framework for large models, aligning closely with the rapid iteration of current models. Our MACEval offers a promising direction that using large models to tackle the evaluation problems for themselves has lower costs than the traditional human-dependent form on broad occasions. Second, the proposed interview-like autonomous evaluation pipeline with open-ended settings is more conducive to exploring the performance boundaries of large models with less subjective bias. Third, although our research is not directly related to the design of generative models, the possibility of spontaneously generating NSFW content in other tasks due to hallucinations or ambiguity in the content of the dialogue cannot be ruled out, since we adopted a human-free autonomous multi-agent collaboration framework in the evaluation process, where the evaluation content is autonomously generated.
Appendix B Detailed Open-Ended Tasks Construction
@@
  • Image Quality Perception (IQP): First, we select a set of keywords (people, animals, nature, plants, cars, objects, buildings, textures, sports, and interiors) based on the categorization of the renowned image website Unsplash [^1] https://unsplash.com/, which served as the source for generating image prompts. Then, we adopt an interviewer-assisted prompt generation strategy to obtain $\mathcal{Q}_{total}$ prompts. Five prevailing text-to-image models, including Stable Diffusion-1.5 (Rombach et al., 2022), Stable-Diffusion-3.5-medium (AI, 2024), Playground v2.5 (Li et al., 2024b), FLUX.1-schnell (Labs, 2024), and Sana (Xie et al., 2024), are randomly invoked externally to generate original image contents. The output image size is set to $512^{2}$ and the num_inference_steps is set to the default value. As for the difficulty control, we generate a set of Gaussian white noise by setting the variance to $1:\infty:1]$ (which represents values starting from $1$ with a step size of $1$) and add it to the original images.
  • Content Understanding (CU): This task evaluates the model’s ability to distinguish image content. We download over 100 different icons from FLATICON [^2] https://www.flaticon.com/ and randomly form part of them into a grid image, which contains smaller icons, as shown in Fig. 8. Then, we examine whether the interviewee model can correctly answer the exact number of target icons in the given grid images. The task difficulty is controlled by the number of grids, i.e., the more grids there are, the more interference terms are presented. We begin with the $3\times 3$ size.
@@
  • Binary Tree Traversal (BTT): Binary tree is a common data structure widely used in computer science and programming. In general, a binary tree structure can be uniquely determined by giving its pre-order and in-order traversals. Based on this, we prompt the interviewer model to generate a binary tree with its pre-order and in-order traversals, then query the interviewee model to output the post-order traversal. The number of nodes is used to control the task difficulty, namely the structural complexity of a binary tree.
@@
  • Code Generation: Similar to the mathematical problems, the difficulty of coding problems is also hard to define. In this paper, we focus on the cyclomatic complexity, which is a software metric that quantifies the complexity of a code’s control flow, specifically measuring the number of linearly independent paths through the code. We first instruct the interviewer model to generate an initial code demand and then send this to the interviewee model for code generation. After that, the generated code will be sent back to the interviewer model for examination. If the interviewee successfully completes the current task, a new requirement will be randomly added, such as incorporating a ‘for’ loop, a ‘while’ loop, or an ‘if’ conditional statement. This process will continue iteratively until the interviewee fails to produce the correct output.
@@
Note that for arithmetic, BTT, and SPS tasks, we implement external code-based supervision to prevent potential generation errors made by the interviewer itself.
@@
Appendix C Additional Information of Evaluation Topology
@@
  • Line: Line-based evaluation topology is the simplest structure where nodes (such as interviewee models and interviewer models) are connected in a sequential manner along a single communication line. In other words, in each round of task evaluation, only one interviewer model engages in dialogue-based assessment with the interviewee model.
@@
Figure 10: Overall performance comparison among 15 LLMs based on the evaluation network.
Appendix D Additional Experimental Results
Subjective Alignment of the Supervisor. We conducted extra subjective experiments where we quantified their alignment with human annotations. Specifically, we adopt two schemes: ’yes-or-no’ evaluation and scoring evaluation (Chen et al., 2024a). 2 experts (with experience in publishing articles on large models) and 8 public participants from campus are invited to participate. We selected four tasks with high error rates (according to Tab. 3), including STUdisrupt., STUmask., Arith.oper., and BTT, for evaluation. We use Claude 3.7 Sonnet and Gemini 2.5 Pro as the supervisor. GPT-4o is the interviewer. Note that in these tasks, we have already implemented external algorithmic function checks for the ground-truth errors (e.g., "23445 + 22784 = ?"). Therefore, human and third-party supervisory models primarily focus on errors related to textual hallucinations. In ‘yes-or-no’ evaluation, the human participants and the supervisor model are instructed to answer “whether there exists any hallucination in the evaluation dialogues? Return yes or no.” In the scoring evaluation, we instruct “Please carefully judge the quality of the evaluation process dialogue based on the task requirements,” and provide a detailed scoring standard with a 1-5 Likert scale:
  • 1: The response is accurate and trustworthy
  • 2: The response is mostly correct but contains minor, non-core factual errors
  • 3: A mix of fact and fabrication makes the response partially unreliable
  • 4: The core claim is false or the response is mostly fabricated and misleading
  • 5: The response is almost entirely invented, nonsensical, or detached from reality.
@@ Appendix D Additional Experimental Results
Table 5: The SRCC between the score of the model and all raters for ’scoring’ setting. The human agreement percentage in the form of experts/crowds for ’yes-or-no’ setting.
@@
Multiple Roles Exploration. To evaluate the effect of different role settings, we further conduct experiments for the same model in interviewee, interviewer, and supervisor. Concretely, we test one open-source model Qwen2.5-7B and one proprietary model GPT-4o in three representative tasks, including STUdisrupt., Arith.scale, and BTT. As shown in Tab. 9, we can observe that changing the supervisor does not explicitly affect the resulting performance. Meanwhile, after changing the interviewer from GPT-4o to Qwen2.5-7B, the evaluated performance of Qwen2.5-7B on STUdisrupt. and Arith.scale tasks grows slightly. We suspect that this is due to the limited diversity of the generated queries stemming from the relatively weak capability of the interviewer model (similar conclusion as Tab. 4). Therefore, we further employ a 1v3 star evaluation topology to solve the mild variance for the same interviewee model across interviewers, which indeed improves the reliability of the evaluation process. Moreover, we find a high error rate (56%) in the BTT task, calculated by the misalignment between the external function and self-generated ground-truth, when Qwen2.5-7B serves as the interviewer. This result is much lower in GPT-4o’s setting, showing the necessity of using relatively powerful models and incorporating external functions.
Table 6: Results on different role settings. ACC-AUC/Max. Perf. is reported.
Interviewee Interviewer Supervisor STUdisrupt. Arith.scale BTT
Qwen2.5-7B GPT-4o Claude 3.7 Sonnet 6.797/8 0.9/1 1.6/2
Qwen2.5-7B Qwen2.5-7B Qwen2.5-7B 6.833/8 1.0/2 1.6/2
Qwen2.5-7B Qwen2.5-7B+GPT-4o+Claude 3.7 Sonnet Qwen2.5-7B 6.802/8 0.9/1 1.6/2
GPT-4o GPT-4o Claude 3.7 Sonnet 8.855/10 1.7/2 3.1/4
GPT-4o GPT-4o GPT-4o 8.855/10 1.7/2 3.1/4
Table 7: Results on different role settings. ACC-AUC/Max. Perf. is reported.
Interviewee Interviewer Supervisor STUdisrupt. Arith.scale BTT
Qwen2.5-7B GPT-4o Claude 3.7 Sonnet 6.797/8 0.9/1 1.6/2
Qwen2.5-7B Qwen2.5-7B Qwen2.5-7B 6.833/8 1.0/2 1.6/2
Qwen2.5-7B Qwen2.5-7B+GPT-4o+Claude 3.7 Sonnet Qwen2.5-7B 6.802/8 0.9/1 1.6/2
GPT-4o GPT-4o Claude 3.7 Sonnet 8.855/10 1.7/2 3.1/4
GPT-4o GPT-4o GPT-4o 8.855/10 1.7/2 3.1/4
@@
  • Cagney can frost a cup cake every <20> seconds and Lacey can frost a cupcake every <30> seconds. Working together, how many cupcakes can they frost in <5> minutes? Avg. Difficulty = 0.134 Avg. Difficulty = 0.134
  • Find the number of pairs of integers (a, b) with <1> $\leqslant a<b\leqslant$ <57> such that $a^{2}$ has a smaller remainder than $b^{2}$ when divided by <57>. Avg. Difficulty = 0.587
  • In a <16>x<16> table of integers, each row and column contains at most <4> distinct integers. What is the maximum number of distinct integers that can be in the whole table? Avg. Difficulty = 0.784
We uniformly sample 10 question templates with incremental difficulty rating from 0.1035 to 0.8548 (The larger the value, the higher the difficulty). During the evaluation data generation, the interviewer model with external function-assisted is instructed to generate query-answer by changing the values in ‘<>’. We set the number of queries at each difficulty level to 100. Rather than relying on manually annotated fine-grained difficulty scores, we increase the number of testing rounds within each predefined difficulty level to obtain a more precise performance estimate. Here, we tested two models: Qwen 2.5-VL-72B and Gemini 1.5 Pro. GPT-4o is used as the interviewer model. In Tab. 8, we observe that the overall declining trend on both dynamic and static settings is similar. The performance under MACEval with multi-round queries exhibits slightly higher values than those with static settings, showing a more genuine performance while demonstrating its potential superiority for longitudinal evaluation against data contamination.
@@
Table 9: Results on different role settings. ACC-AUC/Max. Perf. is reported.
Interviewee Interviewer Supervisor STUdisrupt. Arith.scale BTT
Qwen2.5-7B GPT-4o Claude 3.7 Sonnet 6.797/8 0.9/1 1.6/2
Qwen2.5-7B Qwen2.5-7B Qwen2.5-7B 6.833/8 1.0/2 1.6/2
Qwen2.5-7B Qwen2.5-7B+GPT-4o+Claude 3.7 Sonnet Qwen2.5-7B 6.802/8 0.9/1 1.6/2
GPT-4o GPT-4o Claude 3.7 Sonnet 8.855/10 1.7/2 3.1/4
GPT-4o GPT-4o GPT-4o 8.855/10 1.7/2 3.1/4
#Query STUmask. BTT
10 4.247/5 2.40/3
20 4.256/5 2.50/3
30 4.253/5 2.43/3
50 4.244/5 2.46/3
Table 11: Results of the statistical test.
#Query@N times STUmask. BTT
10@10 [4.2444, 4.2558] [2.4122, 2.4618]
20@10 [4.2457, 4.2533] [2.4265, 2.4570]
p-value 0.8486 0.7227
The Stability of Evaluation Data Generation. We compare the performance differences in terms of the number of generated queries at the same difficulty level on STUmask. and BTT tasks, as shown in Tab. 10. The interviewee and interviewer are Qwen2.5-72B and GPT-4o, respectively. We further conducted two-side t-test and reported 95% PI and p-value for 300 queries in Tab. 11. We can observe that there is no significant difference between the results of 10@10 and 20@10, as the p-value>0.05. Their 95% PIs have a considerable overlap, indicating the robustness of the evaluation process. Furthermore, the 95% PI of 20@10 is more concentrated than that of 10@10, showing the precision of increasing the number of rounds under the given difficulty level.
@@
GPT-4o GPT-4o+Claude 3.7 Sonnet+Gemini 2.5 Pro+DeepSeek V3 55.6/88 41.7/45 58.2/95
@@
Expanding the Task Suite to Include Tasks that are Less Structured and Quantifiable. We further expanded the task suite to include language understanding tasks that are less structured or quantifiable. Specifically, we focus on the following tasks:
  • External material: To avoid data contamination, we manually collected 10 recent news from world economic forum [^3] https://www.weforum.org/ (time period: 2025.7.20-7.25) to ensure that they are not in the training data.
  • Self-generated: The interviewee model is required to generate text contents with a certain length and requirements.
Settings:
  • For task 1-1 (External), where the interviewer randomly requires the interviewee model to paraphrase, simplify, or summarize the given text content with a certain output format, we score tasks purely by their adherence to the instructions (White et al., 2025). The interviewer model is instructed to increase the number of queries in each round until the finishing rate is equal to a very low value. However, in this experiment, we found a relatively slow rate of decline. It is time-consuming to reach the scenario with 0% finishing rate compared to other tasks. Therefore, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries, which can also validate the feasibility of MACEval in handling open language understanding tasks.
  • For task 1-2 (External), we further employ the language bootstrapping (Yang et al., 2024c) (e.g., the proportion of irrelevant words or sentences) to adjust the difficulty rate each success round while satisfying the principle of MACEval (pushing to the limit). Therefore, we report the ACC-AUC and the bootstrapping proportion (%).
  • For task 2 (Self-generated), the interviewer model is instructed to increase the number of queries (difficulty rate) in each round until the finishing rate is equal to a very low value. An external Python function is deployed to count word length for accuracy calculation. Same with scenario 1-1, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries due to the slow decline rate.
  • Four models are selected for evaluation. Three different evaluation topologies are investigated (including 1-hop line and 1v4 star).
In Tab. 12, we find tiny differences in ACC-AUC/Max. Perf. when changing GPT-4o to Claude 3.7 Sonnet under 1-hop line. This can be attributed to the small disparity of the in-process data, which can be mitigated under the 1v4 star evaluation topology, where a mean opinion score (MOS)-based strategy is adopted, showing good robustness and flexibility of the MACEval framework. We notice the results in Task 1-1 and Task-2 are similar, indicating that the influence of the data source in certain tasks is negligible and validating the suitability of AI-generated text in such tasks.
@@
More Evaluation on Math Tasks. We expand the math tasks to solving equations. We choose 8 models for evaluation, where GPT-4o serves as the interviewer model. The difficulty level is set to the number of variables. We require it to retain four decimal places. As listed in Tab. 13, we observe that those open-source models, even with large-scale parameters or mainstream proprietary models, such as GPT-4o and Gemini1.5 Pro, can only solve equations with two variables, showing great potential for general large models to improve. Surprisingly, the cutting-edge Gemini 2.5 Pro, a model skilled in mathematics and reasoning, is only capable of solving systems of three linear equations in three variables to a limited extent.
Table 14: Impact of error propagation.
Error-round 1 Error-round 2 Error-round 3 Error-round 4 Error-round 5
52% 50% 53% 52% 54%
@@
Potential Error Propagating. We conduct extra experiments to explore the effects of errors that are proactively injected into the next round under the current node or evaluation process. Specifically, we add interference characters (unordered and meaningless strings) into the message, which are together taken to the next node. We evaluate the data generation process of the binary tree traversal (BTT) task, which has the highest initial error (52%, Tab. 3). The results in Tab. 14 show that injecting irrelevant textual contents will not affect the inherent error rate that depends on the difficulty of the task itself. Conversely, the progressive accumulation of erroneous content renders the information more readily detectable by a third-party supervisor, facilitating its timely suppression and the issuance of a regeneration directive to the interviewer model. Hence, the applied external algorithm played a crucial role within the single task.
D.1 Comparison to Other Large Model Benchmarks
We compare the evaluation results obtained by MACEval to the popular benchmarks (LiveBench (White et al., 2025) and ChatBot Arena (Chiang et al., 2024)), shown in Fig. 11. It can be observed that the results of MACEval highly correlate with mainstream benchmarks (avg. $r=0.9509$), demonstrating the effectiveness of MACEval and proving the potential of open-ended problem settings.
D.2 Visualization of the Evaluation Stream
To illustrate the working mechanism of MACEval more intuitively, we visualize the dialogue example of the third-party supervisor and prompt templates for both interviewee and interviewer in Fig. 12 and Appendix F. In this paper, we adopt the Claude 3.7 Sonnet (Anthropic, 2024) as the supervisor.

### v2

@@ 5 Conclusion
This work introduces MACEval, a novel dynamic evaluation framework for large generative models, which provides effective solutions to avoid data contamination and reduce human participation by a multi-agent system-based automatic evaluation network. To identify the upper bound of performance and evaluate continually, we propose an AUC-inspired metric, further incorporating the evaluation topologies to measure the overall performance of large models. Experiments on five high-profile capabilities demonstrate the effectiveness, efficiency, and flexibility of MACEval compared to existing benchmarks, presenting a promising direction for achieving fair evaluation.
Impact Statement
Our work holds the potential for significant societal impact, both positive and negative. First, the evaluation is normally considered as important as training in developing large models. An insightful evaluation can offer promising directions to large model developers. The main motivation of this work is to build an automatic, scalable, and trustworthy evaluation framework for large models, aligning closely with the rapid iteration of current models. Our MACEval offers a promising direction that using large models to tackle the evaluation problems for themselves has lower costs than the traditional human-dependent form on broad occasions. Second, the proposed interview-like autonomous evaluation pipeline with open-ended settings is more conducive to exploring the performance boundaries of large models with less subjective bias. Third, although our research is not directly related to the design of generative models, the possibility of spontaneously generating NSFW content in other tasks due to hallucinations or ambiguity in the content of the dialogue cannot be ruled out, since we adopted a human-free autonomous multi-agent collaboration framework in the evaluation process, where the evaluation content is autonomously generated.
Appendix A Detailed Open-Ended Tasks Construction
@@
  • Image Quality Perception (IQP): We select a set of keywords (people, animals, nature, plants, cars, objects, buildings, textures, sports, and interiors) based on the categorization of the renowned image website Unsplash [^1] https://unsplash.com/, which served as the source for generating image prompts. Then, we adopt an interviewer-assisted prompt generation strategy to obtain $\mathcal{Q}_{total}$ prompts. Five prevailing text-to-image models, including Stable Diffusion-1.5 (Rombach et al., 2022), Stable-Diffusion-3.5-medium (AI, 2024), Playground v2.5 (Li et al., 2024b), FLUX.1-schnell (Labs, 2024), and Sana (Xie et al., 2024), are randomly invoked externally to generate original image contents. The output image size is set to $512^{2}$, and the num_inference_steps is set to the default value. As for the difficulty control, we follow the adjustable parameters and ranges in Image Adjustment Tool Palette (Canon, 2024) and generate a set of Gaussian white noise by setting the variance to $1:\infty:1]$ (which represents values starting from $1$ with a step size of $1$) and add it to the original images.
  • Content Understanding (CU): This task evaluates the model’s ability to distinguish image content. We download over 100 different source icons from FLATICON [^2] https://www.flaticon.com/ and randomly form part of them into a grid image, which explicitly avoids image content contamination, as shown in Fig. 10. Then, we examine whether the interviewee model can correctly answer the exact number of target icons in the given grid images. The task difficulty is controlled by the number of grids, i.e., the more grids there are, the more interference terms are presented. We begin with the $3\times 3$ size.
@@
  • Binary Tree Traversal (BTT): The binary tree is a common data structure widely used in computer science and programming. In general, a binary tree structure can be uniquely determined by given its pre-order and in-order traversals. Based on this, we prompt the interviewer model to generate a binary tree with its pre-order and in-order traversals, then query the interviewee model to output the post-order traversal. The number of nodes is used to control the task difficulty, namely, the structural complexity of a binary tree.
@@
  • Code Generation: Similar to the mathematical problems, the difficulty of coding problems is also hard to define. In this paper, we focus on the cyclomatic complexity, which is a software metric that quantifies the complexity of a code’s control flow, specifically measuring the number of linearly independent paths through the code. We first instruct the interviewer model to generate an initial code demand and then send this to the interviewee model for code generation. After that, the generated code will be sent back to the interviewer model for examination. If the interviewee completes the current task, a new requirement will be randomly added, such as incorporating a ‘for’ loop, a ‘while’ loop, or an ‘if’ conditional statement. This process will continue iteratively until the interviewee fails to produce the correct output.
@@
Note that for arithmetic, BTT, and SPS tasks, we implement external code-based supervision to prevent potential generation errors made by the interviewers themselves, shown in Fig. 12.
@@
Figure 12: External upervision functions used in the BTT and SPS tasks.
Appendix B Additional Information of Evaluation Topology
@@
  • Line: Line-based evaluation topology is the simplest structure where nodes (such as interviewee models and interviewer models) are connected sequentially along a single communication line. In other words, in each round of task evaluation, only one interviewer model engages in dialogue-based assessment with the interviewee model.
@@
Figure 13: Overall performance comparison among 15 LLMs based on the right evaluation network, where colored clusters represent evaluation agents for different capabilities and red arrows denote activated evaluation routes.
Appendix C Additional Experimental Results
Subjective Alignment of the Supervisor. We conducted extra subjective experiments where we quantified their alignment with human annotations. Specifically, we adopt two schemes: ‘yes-or-no’ evaluation and scoring evaluation (Chen et al., 2024a). 2 experts (with experience in publishing articles on large models) and 8 public participants from the campus are invited to participate. We selected four tasks with high error rates (according to Tab. 4), including STUdisrupt., STUmask., Arith.oper., and BTT, for evaluation. We use Claude 3.7 Sonnet and Gemini 2.5 Pro as the supervisor. GPT-4o is the interviewer. Note that in these tasks, we have already implemented external algorithmic function checks for the ground-truth errors (e.g., “23445 + 22784 = ?”). Therefore, human and third-party supervisory models primarily focus on errors related to textual hallucinations. In ‘yes-or-no’ evaluation, the human participants and the supervisor model are instructed to answer “whether there exists any hallucination in the evaluation dialogues? Return yes or no.” In the scoring evaluation, we instruct “Please carefully judge the quality of the evaluation process dialogue based on the task requirements,” and provide a detailed scoring standard with a 1-5 Likert scale:
  • Level-1: The response is accurate and trustworthy
  • Level-2: The response is mostly correct but contains minor, non-core factual errors
  • Level-3: A mix of fact and fabrication makes the response partially unreliable
  • Level-4: The core claim is false, or the response is mostly fabricated and misleading
  • Level-5: The response is almost entirely invented, nonsensical, or detached from reality.
@@ Appendix C Additional Experimental Results
Table 7: The Spearman’s $r$ between the score of the model and all raters for ‘scoring’ setting. The human agreement percentage in the form of experts/crowds for a ‘yes-or-no’ setting.
@@
  • Cagney can frost a cup cake every <20><20> seconds and Lacey can frost a cupcake every <30><30> seconds. Working together, how many cupcakes can they frost in <5><5> minutes?—Avg. Difficulty = 0.134
  • Find the number of pairs of integers (a, b) with <1><1> $\leqslant a<b\leqslant$ <57><57> such that $a^{2}$ has a smaller remainder than $b^{2}$ when divided by <57><57>.—Avg. Difficulty = 0.587
  • In a $16>\times<16>$ table of integers, each row and column contains at most <4><4> distinct integers. What is the maximum number of distinct integers that can be in the whole table?—Avg. Difficulty = 0.784
We uniformly sample 10 question templates with incremental difficulty rating from 0.1035 to 0.8548 (The larger the value, the higher the difficulty). During the evaluation data generation, the interviewer model with an external function-assisted is instructed to generate a query-answer by changing the values in ‘<><>’. We set the number of queries at each difficulty level to 100. Rather than relying on manually annotated fine-grained difficulty scores, we increase the number of testing rounds within each predefined difficulty level to obtain a more precise performance estimate. Here, we tested two models: Qwen 2.5-VL-72B and Gemini 1.5 Pro. GPT-4o is used as the interviewer model. In Tab. 8, we observe that the overall declining trend on both dynamic and static settings is similar. The performance under MACEval with multi-round queries exhibits slightly higher values than those with static settings, showing a more genuine performance while demonstrating its potential superiority for longitudinal evaluation against data contamination.
@@
GPT-4o GPT-4o, Claude 3.7 Sonnet, Gemini 2.5 Pro, DeepSeek V3 55.6/88 41.7/45 58.2/95
@@
Results on Tasks that are Less Structured and Quantifiable. We further expanded the task suite to include language understanding tasks that are less structured or quantifiable. Specifically, we focus on the following tasks. 1) External material: To avoid data contamination, we manually collected 10 recent news articles from world economic forum [^3] https://www.weforum.org/ (time period: 2025.7.20-7.25, later than the cutoff date of the evaluated models) to ensure that they are not in the training data. 2) Self-generated: The interviewee model is required to generate text contents with a certain length and requirements. Below we provide detailed settings.
  • Task 1-1 (External): The interviewer randomly requires the interviewee model to paraphrase, simplify, or summarize the given text content with a certain output format. We score tasks purely by their adherence to the instructions (White et al., 2025). The interviewer models are instructed to increase the number of queries in each round until the finishing rate equals a very low value. However, in these experiments, we found a relatively slow rate of decline. It is time-consuming to reach the scenario with 0% finishing rate compared to other tasks. Therefore, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries, which can also validate the feasibility of MACEval in handling open language understanding tasks.
  • Task 1-2 (External): We further employ the language bootstrapping (Yang et al., 2024c) (e.g., the proportion of irrelevant words or sentences) to adjust the difficulty rate each success round while satisfying the principle of MACEval (pushing to the limit). Therefore, we report the ACC-AUC and the bootstrapping proportion (%).
  • Task 2 (Self-generated): The interviewer models are instructed to increase the number of queries (difficulty rate) in each round until the finishing rate equals a very low value. An external Python function is deployed to count word length for accuracy calculation. Same with scenario 1-1, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries due to the slow decline rate.
As shown in Tab. 9, four models are selected for evaluation, and three different evaluation topologies are investigated (including 1-hop line and 1v4 star). We find a tiny difference in ACC-AUC/Max. Perf. when changing GPT-4o to Claude 3.7 Sonnet under 1-hop line. This can be attributed to the small disparity of the in-process data, which can be mitigated under the 1v4 star evaluation topology, where a mean opinion score (MOS)-based strategy is adopted, showing good robustness and flexibility of the MACEval framework. We notice the results in Task 1-1 and Task-2 are similar, indicating that the influence of the data source in certain tasks is negligible and validating the suitability of AI-generated text in such tasks.
More Evaluation on Math Tasks. We expand the math tasks to solving the equations, and choose 8 models for evaluation, where GPT-4o serves as the interviewer model. The difficulty level is set to the number of variables. We require it to retain four decimal places. As listed in Tab. 10, we observe that those open-source models, even with large-scale parameters or mainstream proprietary models, such as GPT-4o and Gemini1.5 Pro, can only solve equations with two variables, showing great potential for general large models to improve. Surprisingly, the cutting-edge Gemini 2.5 Pro, a model skilled in mathematics and reasoning, is only capable of solving systems of three linear equations in three variables to a limited extent.
@@
C.1 Visualization of the Evaluation Stream
To illustrate the working mechanism of MACEval more intuitively, we visualize the dialogue example of the third-party supervisor and prompt templates for both interviewee and interviewer in Fig. 14 and Appendix G. In this paper, we adopt the Claude 3.7 Sonnet (Anthropic, 2024) as the supervisor.
Appendix D Limitations
There are two potential limitations acknowledged below. First, we apply evaluation topologies with fewer hops, i.e., nodes and edges, to preliminarily demonstrate the effectiveness of MACEval. In the future, it is possible to construct larger evaluation networks to include a richer array of capability assessments. Second, in MACEval, the evaluation data stream is composed of question-answer pairs graded by difficulty. One of the principles in constructing such open-ended tasks is that the difficulty should be quantifiable, which imposes limitations on the implementation of many tasks.